OpenStudy (morgengarner):

Simplify the expression. The quantity of the quantity 3 multiplied by x end of quantity divided by y end of quantity raised to the power of negative 3. A. 27x3 y3 B. Y raised to the power of 3 divided by the quantity 27 multiplied by x raised to the power of 3 end of quantity. C. 27 divided by the quantity x raised to the power of 3 multiplied by y raised to the power of 3 end of quantity. D. The quantity 27 multiplied by x raised to the power of 3 end of quantity divided by y raised to the power of 3.

OpenStudy (johnweldon1993):

\[\large (\frac{3x}{y})^3\] Remember \(\large (a^x)^y =a^{xy}\) and \[\large (\frac{a}{b})^z = \frac{a^z}{b^z}\] So here \[\large (\frac{3x}{y})^3 = \frac{(3x)^3}{y^3} = \frac{3^3x^3}{y^3}= ?\]